# First Statistical Detection of Cool Gas Outflows with JWST Towards Cosmic Dawn

### v1

@@ 3.2 Outflow Properties
The boxcar method, while model-independent, can yield non-physical results under specific conditions. In the typical observed outflows, one expects $\rm EW_{out}>0$ and $v_{\rm out}<0$. However, in the presence of significant noise or complex kinematics, the method may produce $\rm EW_{out}<0$ or $v_{\rm out}>0$. We note that $\rm EW_{out}<0$ indicates more absorption in the red wing, which is suggestive of inflow features. However, this method tailored for outflow characterization is not suitable for the corresponding analysis. When $\rm EW_{out}$ is small, the derived $v_{\rm out}$ becomes unstable to and may diverge to spuriously large values, which should be treated with caution.
To validate the robustness of our stacking analysis, we also implement the same boxcar measurement on individual galaxy spectra. First we identify potential outflow candidates by conducting a visual inspection of all JWST spectra which covers Mg II region, selecting those with discernible Mg II absorption features. For these candidates, we calculate the Mg II absorption EW by integrating the absorbed flux between $2790\ \rm\AA$ and $2808\ \rm\AA$ . The uncertainty in the EW, $\rm EW\_err$, is derived from the error of the flux densities. We then apply a S/N cut of $\rm EW\_err>5$, which yields a final sample of 42 individual JWST spectra with significant Mg II absorption. The properties and spectra of these sources are tabulated and presented in the Appendix B.
4 Results and Discussion
4.1 Absorption Feature Across Cosmic Time
@@
Figure 2 shows the final stacked spectra for five distinct redshift bins. Our S/N-weighted stacking technique successfully reveals clear absorption features for both the Mg II and Na I D doublets across the cosmic timeline probed by our sample ($1<z<10$). Despite showing greater spectral fluctuations, the stacked spectra produced using equal-weight stacking also exhibit consistent blueshifted absorption features (see left panels of Figure 7). This consistency confirms that the detected outflow signal is robust and not an artifact of our weighting scheme. The left panels reveal consistently blueshifted absorption troughs relative to the systemic velocities (marked by the red dashed lines), providing clear evidence for the prevalent cool outflows. The S/N of the composite spectra varies across redshift bins, a difference primarily attributed to the number and intrinsic quality of the spectra contributing to each stack. We note that here we focus on establishing the detection of these blueshifted outflows across cosmic time. A robust quantification of their evolutionary trend is not feasible with the current dataset, as the underlying distributions of key galaxy properties (e.g., mass, star formation rate) differ between redshift bins.
The Na I D doublet, shown in the right panels, is also detected in the stacked spectra, showing blueshifted absorption features. However, as shown in Figure 2, the blue wing of the absorption is blended with the strong He I $\lambda 5876$ emission line, hindering local continuum subtraction. Furthermore, the Na I D absorption is less prominent than Mg II.
More fundamentally, the observed Na I D absorption is a composite signal. It originates not only from cool, neutral gas in the ISM—the component tracing the outflow—but also from the stellar photospheres within the host galaxy’s stellar population. This stellar absorption component arises predominantly from the atmospheres of cool, evolved stars (e.g., K and M giants and supergiants), which can contribute substantially to the total observed EW (e.g., Heckman2000; Rupke2005). Disentangling the ISM contribution from this stellar contamination is highly non-trivial and requires sophisticated full spectral fitting with stellar population synthesis (SPS) models. Such modeling is fraught with uncertainties, especially for high-redshift galaxies where the stellar populations are less well-constrained.
@@
4.2 Absorption Feature Across Stellar Mass Bins
Figure 3: Stacked Mg II feature in stellar mass and redshift bins. Each colored box corresponds to a specific redshift bin, as indicated by the legend on the right. Within each box, the panels are arranged in order of increasing stellar mass from top to bottom. The stellar mass range and the number of spectra (NN) in each stack are annotated in each panel. The blue solid line, shaded uncertainty region, and dashed lines are the same as in Figure 2.
@@ 4.2 Absorption Feature Across Stellar Mass Bins
However, the outflow signal in the $1<z<2$ bins is notably not significant, yielding small $\rm EW_{out}$ values and unstable, spuriously large $v_{\rm out}$ values. This is likely attributable to the limited sample size and the low spectral S/N of JWST galaxies with Mg II coverage in this redshift range, despite previous studies having robustly detected outflows at similar epochs (Weiner2009; Bordoloi2014; Yu2025). In contrast, the $2<z<3$ and $3<z<4$ bins exhibit high S/N composite spectra, enabling well-constrained measurements of $\rm EW_{out}$ and $v_{\rm out}$ suitable for statistical analysis. For the highest redshift bins ($4<z<5$ and $5<z<10$), the absorption profiles, though noisy, display clear blueshifts. This provides strong evidence that cool gas outflows are already present at $z>5$ (corresponding $\sim$ 1 Gyr of the Universe) galaxies. To our knowledge, this is the first statistical detection of outflow signatures in $z>3$ galaxies, indicating that feedback mechanisms plays a critical role in the early phase of galaxy evolution.
Figure 4: Scaling relations for the Mg II absorption EW. Left panel: The outflow EW as a function of stellar mass. Our stacked measurements are shown as star symbols, color-coded by their weighted mean redshift. The grey dashed line shows a linear fit to the stacked data points. For comparison, small circles show the measurements from the individual galaxies presented in the Appendix B, color-coded by their redshift. Right panel: The outflow EW as a function of redshift, with an additional top axis showing cosmic time. The star symbols again represent our stacked results, but all data points are now color-coded by their weighted mean stellar mass, as indicated by the colorbar. The grey dashed line is a linear fit to the stacked data.
Figure 5: The outflow velocity versus stellar mass relation. Our measurements from the weighted stacking are shown as star symbols. The color of our data points corresponds to the mean redshift of the stack, as indicated by the color bar on the right. We compare our results with literature data from Weiner2009, Bordoloi2014, Davis2023, Yu2025, Liboni2025, and Valentino2025, as detailed in the legend. Our work provides new constraints, particularly at the low-mass, high-redshift ($z>3$) end of the parameter space.
4.3 Scaling Relations of Outflow Properties
The scaling relations for the measured Mg II $\rm EW_{out}$ are displayed in Figure 4. Our stacked measurements (star symbols) demonstrate that EWout systematically increases with stellar mass across all cosmic epochs. A linear fit to the stacked data (grey dashed line) confirms this dominant trend, and we find no significant offset for the high-redshift stacks, indicating that galaxies from all epochs broadly follow a universal mass-$\rm EW_{out}$ relation. For comparison, we also plot individual detections from our gallery in Appendix B (small circles), which, despite exhibiting significant scatter at a fixed mass, follow the same average trend. These sources systematically exhibit larger $\rm EW_{out}$ than the stacked results, indicating that selection effects plays a role in analyzing individual detections.
The right panel of Figure 4 shows the outflow EW as a function of redshift. No simple, monotonic trend with cosmic time is detected, which is primarily a consequence of the strong mass-dependence of the EW, combined with the fact that the average stellar mass of our stacked samples evolves with redshift (as visualized by the color-coding of the star symbols). For instance, the stacked point at $z\sim 6$ has a lower EW than the point at $z\sim 3$, not necessarily because outflows are intrinsically weaker, but because the average mass of the $z\sim 6$ stack is significantly lower. Taken together, these two panels strongly indicate that stellar mass of galaxies is linked to the cool gas outflow EW.
A comparison between our measured outflow velocities and a compilation of literature results is presented in Figure 5. Our measurements (star symbols) significantly extend previous works into the low-mass, high-redshift frontier. A key result from this comparison is the remarkable consistency of outflow velocities at $z>3$. Across our redshift bins from $z\sim 3$ up to $z\sim 10$, the outflow velocities remain robustly centered around $\sim 350\,\mathrm{km\,s^{-1}}$ and do not exhibit significant evolution. This lack of evolution at high redshifts is a primary finding of our work.
For context, cool outflow velocities in the literature vary with galaxy properties. For instance, vigorous starbursts in merging galaxies can drive outflows up to $\sim 1000\,\mathrm{km\,s^{-1}}$ (Davis2023), whereas quiescent galaxies exhibit slower winds (Valentino2025). Our results for typical star-forming galaxies at $z>3$ establish a crucial baseline in a previously unexplored regime. While there might be an evolution in outflow properties from $z<2$ to $z>2$, our data robustly show that for $z>3$, the relationship between outflow velocity and stellar mass is largely unevolving.
Furthermore, we find that the cool outflow velocities in the literature varies with the properties of galaxies. For instance, vigorous starbursts in merging galaxies can drive outflows up to $\sim 1000~\rm km,s^{-1}$ (Davis2023), whereas quiescent galaxies exhibit slower winds (Valentino2025). The mean outflow velocity measured from our individual spectra (mostly at $3<z<4$) are consistent with the scaling relation of Yu2025, which was established using low-redshift star-forming main sequence (SFMS) galaxies. Nonetheless, our stacked results yield systematically higher velocities, particularly at the low-mass end. Although we have claimed that the measurement in the low-mass bin of $1<z<2$ sample could be affected by numerical instability, the elevated velocities in higher-redshift low-mass bins appear robust. This indicates that the high-redshift JWST galaxies in our study may be experiencing more intense, bursty star formation, potentially driving stronger outflows (Looser2025; Clarke2025; Lyu2025). This also suggests that the dynamics of the cool gas outflows before the cosmic noon may differ fundamentally from those in the low-redshift universe.
Crucially, thanks to the large and deep JWST sample, we provide statistical constraints on the velocities of cool gas outflows in low-mass ($M_{*}\lesssim 10^{9.5}~\rm{M_{\odot}}$) galaxies at $z>3$. These measurements in the low-mass, high-redshift regime are vital for constraining feedback models, as they probe the numerous progenitors of today’s typical galaxies during their most active growth phases. While there is some evidence for an evolution in the $v_{\rm out}-M_{*}$ relation from $z<1$ (literature) to $z\approx 2$, we find no significant further evolution at higher redshifts ($z>3$).
Figure 6 presents another view of our results, showing the outflow velocity as a function of cosmic time. At first glance, the evolution of our measured outflow velocities (star symbols) with redshift appears complex, with no simple monotonic trend. However, this is primarily a consequence of selection effects, as the mean stellar mass of our stacked samples generally decreases with increasing redshift, which is clearly visualized by the color-coding of the data points. This figure starkly illustrates the unique contribution of our work: providing the statistically-derived outflow velocity measurements for typical star-forming galaxies at $z>3$ and on the approach to the Cosmic Dawn.
Combining the results from Figures 4, 5, and 6, This lack of significant evolution in the outflow properties at $z>3$ provides a powerful test for theoretical models of galaxy feedback in the early universe. For instance, the feedback-free starburst (FFB) model (Dekel2023; Li2024) predicts a fundamental transition in the mode of star formation for massive galaxies at $z\gtrsim 8-10$, where feedback is posited to become inefficient. Such a dramatic change in the interplay between star formation and feedback would presumably lead to a significant alteration in outflow properties, or perhaps even a suppression of the strong, organized outflows that we observe. Contrary to this prediction, our findings reveal a remarkable consistency in outflow properties from $z\sim 3$ up to $z\sim 10$. The fact that galaxies at $z>5$ follow similar scaling relations and exhibit similar outflow velocities to those at $z\sim 3-4$ suggests a persistent, unevolving feedback mechanism is already in place by $z=10$. This result challenges the scenario where the FFB mechanism is the dominant mode governing the baryon cycle in the typical star-forming galaxies probed by our sample, placing a strong constraint on its applicability.
Figure 6: The cosmic evolution of outflow velocity. Our measurements from the full redshift bins stacks are shown as large star symbols. The color of all data points corresponds to the weighted mean stellar mass of the respective galaxy sample, as indicated by the color bar on the right. For context, we compare our results with a compilation of literature data from Weiner2009, Bordoloi2014, Davis2023, Yu2025, Liboni2025, and Valentino2025.
Table 1: Properties of Stacked Samples and Outflow Measurements. The columns are as follows: (1) Redshift bin. (2) Weighted mean redshift of the bin. (3) Stellar mass bin. For rows with a dash, measurements correspond to the entire redshift bin. (4) Weighted mean of the logarithm of the stellar mass. (5) Number of spectra in the stack. (6) Weighted mean outflow velocity derived from the Mg II doublet. (7) Outflow equivalent width of the Mg II doublet.
$z$ $\langle z\rangle_{w}$ $\log M_{*}$ $\langle\log M_{*}\rangle_{w}$ NN $v_{\rm out}$ $\rm{EW_{\rm out}}$
($\rm M_{\odot}$) ($\rm M_{\odot}$) ($\rm km\,s^{-1}$) (Å)
1–2 1.836 9.65 334 $480^{+394}_{-455}$ $0.54^{+0.70}_{-0.68}$
1–2 1.828 8.0–9.5 8.96 224 $370^{+802}_{-590}$ $0.20^{+1.38}_{-1.23}$
1–2 1.842 9.5–12.0 10.14 110 $381^{+106}_{-183}$ $0.83^{+0.77}_{-0.88}$
2–3 2.602 9.47 763 $364^{+32}_{-31}$ $1.75^{+0.38}_{-0.36}$
2–3 2.608 8.0–9.5 8.99 527 $338^{+104}_{-112}$ $0.96^{+0.56}_{-0.49}$
2–3 2.582 9.5–10.5 9.93 175 $363^{+33}_{-30}$ $2.21^{+0.42}_{-0.56}$
2–3 2.648 10.5–12.0 10.78 61 $350^{+33}_{-45}$ $3.92^{+1.38}_{-1.24}$
3–4 3.458 9.39 586 $260^{+23}_{-21}$ $2.02^{+0.32}_{-0.30}$
3–4 3.486 8.0–9.5 8.95 428 $264^{+53}_{-43}$ $1.44^{+0.43}_{-0.43}$
3–4 3.410 9.5–10.5 9.88 124 $248^{+40}_{-27}$ $1.89^{+0.54}_{-0.53}$
3–4 3.423 10.5–12.0 10.88 34 $314^{+39}_{-28}$ $4.75^{+0.96}_{-1.15}$
4–5 4.356 9.24 210 $399^{+84}_{-189}$ $1.26^{+0.61}_{-0.74}$
4–5 4.385 8.0–9.5 8.89 174 $357^{+108}_{-163}$ $1.23^{+0.69}_{-0.75}$
4–5 4.281 9.5–12.0 10.14 36 $380^{+240}_{-255}$ $1.30^{+1.32}_{-1.44}$
5–10 6.067 9.01 497 $336^{+57}_{-85}$ $1.81^{+0.75}_{-0.70}$
5–10 6.078 8.0–9.5 8.81 450 $323^{+97}_{-98}$ $1.23^{+0.76}_{-0.71}$
5–10 6.017 9.5–12.0 9.84 47 $335^{+65}_{-91}$ $3.18^{+1.62}_{-1.85}$

### v2

@@ III.2 Outflow Properties
In this analysis, the absorption profiles of the doublet are assumed to be identical. This assumption is supported by observational evidence that the doublet ratio in most Mg II absorbers is close to 1:1 due to saturation, indicating that the cool, neutral phase gas traced by Mg II is clumpy and optically thick (e.g., Lan and Fukugita, 2017). Accordingly, the mean velocity of a single absorption trough is derived by merging the spectrum blueward of Mg II $\lambda$2796 with that redward of Mg II $\lambda$2803 and then computing the EW-weighted mean offset in the velocity space.
@@
To validate the robustness of our stacking analysis, we also implement the same boxcar measurement on individual galaxy spectra. First we identify potential outflow candidates by conducting a visual inspection of all JWST spectra which covers Mg II region, selecting those with discernible Mg II absorption features. For these candidates, we calculate the Mg II absorption EW by integrating the absorbed flux between $2790\ \rm\AA$ and $2808\ \rm\AA$ . The uncertainty in the EW, $\rm EW_{\rm err}$, is derived from the error of the flux densities. We then apply a S/N cut of $\rm EW_{\rm err}>5$, which yields a final sample of 42 individual JWST spectra with significant Mg II absorption. The properties and spectra of these sources are tabulated and presented in the Appendix B.
IV Results
IV.1 Absorption Feature Across Cosmic Time
[…]
Figure 2 shows the final stacked spectra for five distinct redshift bins. Our S/N-weighted stacking technique successfully reveals clear absorption features for both the Mg II and Na I D doublets across the cosmic timeline probed by our sample ($1<z<10$). For Mg II doublets, despite showing greater spectral fluctuations, the stacked spectra produced using equal-weight stacking exhibit consistent blueshifted absorption features relative to the systemic velocities (marked by the red dashed lines in the left panels of Figure 6). This consistency confirms that the detected outflow signal is robust and not an artifact of our weighting scheme. The S/N of the composite spectra varies across redshift bins, a difference primarily attributed to the number and intrinsic quality of the spectra contributing to each stack. We note that here we focus on establishing the detection of these blueshifted outflows across cosmic time. A robust quantification of their evolutionary trend is not feasible with the current dataset, as the underlying distributions of key galaxy properties (e.g., mass, star formation rate) differ between redshift bins.
The Na I D doublet, shown in the right panels, is also detected in the stacked spectra. However, unlike the Mg II doublet, the Na I D profiles do not exhibit clear blueshifted features. As shown in Figure 2, the blue wing of the absorption (where outflow signatures are expected) is severely blended with the strong He I $\lambda$5876 emission line, hindering both local continuum subtraction and kinematic characterization.
More fundamentally, the observed Na I D absorption is a composite signal. It originates not only from cool, neutral gas in the ISM—the component tracing the outflow—but also from the stellar photospheres within the host galaxy’s stellar population. This stellar absorption component arises predominantly from the atmospheres of cool, evolved stars (e.g., K and M giants and supergiants), which can contribute substantially to the total observed EW (e.g., Heckman et al., 2000; Rupke et al., 2005). Disentangling the ISM contribution from this stellar contamination is highly non-trivial and requires sophisticated full spectral fitting with stellar population synthesis (SPS) models. Such modeling is fraught with uncertainties, especially for high-redshift galaxies where the stellar populations are less well-constrained.
@@ IV.1 Absorption Feature Across Cosmic Time
IV.2 Absorption Feature Across Stellar Mass Bins
Figure 3: Stacked Mg II feature in stellar mass and redshift bins. Each colored box corresponds to a specific redshift bin, as indicated by the legend on the right. Within each box, the panels are arranged in order of increasing stellar mass from top to bottom. The stellar mass range and the number of spectra in each stack are annotated in each panel. The blue solid line, shaded uncertainty region, and dashed lines are the same as in Figure 2.
@@ IV.2 Absorption Feature Across Stellar Mass Bins
However, the outflow signal in the $1<z<2$ bins is notably not significant, yielding small $\rm EW_{out}$ values and unstable, spuriously large $v_{\rm out}$ values. This is likely attributable to the limited sample size and the low spectral S/N of JWST galaxies with Mg II coverage in this redshift range, despite previous studies having robustly detected outflows at similar epochs (Weiner et al., 2009; Bordoloi et al., 2014; Yu et al., 2025). In contrast, the $2<z<3$ and $3<z<4$ bins exhibit high S/N composite spectra, enabling well-constrained measurements of $\rm EW_{out}$ and $v_{\rm out}$ suitable for statistical analysis. For the highest redshift bins ($4<z<5$ and $5<z<10$), the absorption profiles, though noisy, display clear blueshifts. This provides strong evidence that cool gas outflows are already present at $z>5$ (corresponding $\sim$ 1 Gyr of the Universe) galaxies. Furthermore, this is the first statistical detection of Mg II-traced outflow signatures in $z>3$ galaxies. This finding complements recent studies of ionized outflows (e.g., Carniani et al., 2024; Saldana-Lopez et al., 2025; Cooper et al., 2025) by confirming that powerful feedback mechanisms are not only present but also effective at entraining the cool gas during the early stages of galaxy evolution.
@@
Figure 5: The outflow velocity versus stellar mass relation. Our measurements from the weighted stacking are shown as star symbols. The color of our data points corresponds to the weighted mean redshift of the stack, as indicated by the color bar on the right. We compare our results with literature data from Weiner et al. (2009), Bordoloi et al. (2014),Sugahara et al. (2019), Davis et al. (2023), Yu et al. (2025), Liboni et al. (2025), Valentino et al. (2025),Cooper et al. (2025), Wu (2025) and Xu et al. (2025), as detailed in the legend. The colored solid curves represent the estimated escape velocities ($v_{\rm esc}\approx 3v_{\rm circ}$) as a function of stellar mass at different redshifts ($z=1$ to $z=6$, color-coded), derived using the SHMR of Moster et al. (2013).
IV.3 Scaling Relations of Absorption Equivalent Width
The scaling relations for the measured Mg II $\rm EW_{out}$ are displayed in Figure 4. Our stacked measurements (star symbols) demonstrate that EWout systematically increases with stellar mass across all cosmic epochs. To quantify these trends, we perform a weighted linear least-squares fit to the observed stacked data points $(x_{i},y_{i}\pm\sigma_{i})$ using the functional form of $y=ax+b$, where $a$ is the slope, $b$ is the intercept, and $\sigma_{i}$ denotes the measurement uncertainty (the mean of the upper and lower errors) of the $i$-th data point. The best-fit parameters are estimated by minimizing the weighted residual sum-of-squares $\chi^{2}=\sum_{i}[(y_{i}-ax_{i}-b)/\sigma_{i}]^{2}$ via the scipy.optimize.curve_fit algorithm (Virtanen et al., 2020). To ensure the robustness of our results, we exclude the stacked data points at $1<z<2$ from the fitting process (for all scaling relations in this work). For the relation between outflow equivalent width and stellar mass of stacked points (Figure 4, left panel), the derived fitting relation is:
$\rm EW_{out}=(1.21\pm 0.35)\log(M_{*}/M_{\odot})+(-9.68\pm 3.29).$ (6)
The measurements for the $z>3$ bins (including the highest-redshift $5<z<10$ bin) are all consistent with the best-fit relation within their $1\sigma$ measurement uncertainties, indicating that galaxies from all epochs broadly follow a universal mass-$\rm EW_{out}$ relation. For comparison, we also plot individual detections from our gallery in Appendix B (small circles), which, despite exhibiting significant scatter at a fixed mass, follow the same average trend. These sources systematically exhibit larger $\rm EW_{out}$ than the stacked results, indicating that selection effects plays a role in analyzing individual detections.
The right panel of Figure 4 shows the outflow EW as a function of redshift. The slope of the redshift relation (in the legend) is statistically consistent with zero, indicating the absence of a simple, monotonic trend with cosmic time. This is primarily a consequence of the strong mass-dependence of the EW, combined with the evolution of the average stellar mass across our stacked samples (as visualized by the color-coding). For instance, the stacked point at $z\sim 6$ has a lower EW than the point at $z\sim 3$, not necessarily because outflows are intrinsically weaker, but because the average mass of the $z\sim 6$ stack is significantly lower. Taken together, these two panels strongly indicate that stellar mass of galaxies is linked to the cool gas outflow EW.
IV.4 Outflow Velocities
Figure 5 presents our measured outflow velocities as a function of host stellar mass, compared with a compilation of literature results. Our measurements (star symbols) significantly extend previous works of cool gas outflows into the low-mass, high-redshift frontier. A key result from this comparison is the remarkable consistency of outflow velocities at $z>3$. Across our redshift bins from $z\sim 3$ up to $z\sim 6$ (weighted mean redshift of the highest redshift bin), the outflow velocities remain robustly centered around $\sim 300-400\,\mathrm{km\,s^{-1}}$ and do not exhibit significant evolution.
Cool outflow velocities reported in the literature vary significantly with galaxy properties. For instance, vigorous starbursts in merging galaxies can drive outflows with average velocities reaching up to $\sim 1000~\rm km\,s^{-1}$ (Davis et al., 2023). In contrast, quiescent galaxies typically exhibit significantly slower winds (Wu, 2025; Valentino et al., 2025). This difference may stem from the depletion of gas reservoirs and the lack of fresh gas replenishment in quenched systems, whereas typical star-forming galaxies are continuously sustained by gas accretion that fuels energetic feedback. We compare with consistent methodologies from the literature: Weiner et al. (2009) co-added over a thousand spectra and decomposed the Mg II absorption trough into symmetric and outflow components (we plot the median velocity of their outflow component); Bordoloi et al. (2014) employed a boxcar integration method on stacked Mg II spectra, which is methodologically similar to this work. Our stacked results for typical star-forming galaxies at $z>3$ establish a crucial baseline in this high-redshift regime. While there may be an evolution in outflow properties from $z<2$ to $z>2$, our data reveal no evidence for evolution in the scaling relation between outflow velocity and stellar mass at $z>3$.
The mean outflow velocity measured from our individual spectra (mostly at $3<z<4$) are consistent with the scaling relation of Yu et al. (2025), which was established by stacking low-redshift star-forming main sequence (SFMS) galaxies. Nonetheless, our stacked results yield systematically higher velocities, particularly at the low-mass end. Although we have claimed that the measurement in the low-mass bin of $1<z<2$ sample could be affected by numerical instability, the elevated velocities in higher-redshift low-mass bins appear robust. This indicates that the high-redshift JWST galaxies in our study may be experiencing more intense, bursty star formation, potentially driving stronger outflows (Looser et al., 2025; Clarke et al., 2025; Lyu et al., 2025a). This also suggests that the dynamics of the cool gas outflows before the cosmic noon may differ fundamentally from those in the low-redshift universe.
Crucially, we provide statistical constraints on the velocities of cool gas outflows in low-mass ($M_{*}\lesssim 10^{9.5}~\rm{M_{\odot}}$) galaxies at $z>3$ and towards cosmic dawn. These measurements are vital for constraining feedback models, as they probe the numerous progenitors of today’s typical galaxies during their most active growth phases. While there is some evidence for an evolution in the $v_{\rm out}-M_{*}$ relation from $z<1$ (literature) to $z\approx 2$, we find no significant further evolution at higher redshifts ($z>3$). We further examine the outflow velocity as a function of cosmic time in Appendix C (Figure 8).
Recently, Glazer et al. (2025) reported a small velocity offset for low-ionization species (LIS) absorption ($v_{\rm cen,LIS}=-23.5\pm 50.7\,\rm km\,s^{-1}$) from their unweighted mean stack of $z\sim 7$ galaxies, in tension with our measured outflow velocities at similar redshifts. However, this discrepancy primarily stems from fundamental differences in methodology. Our work uses a S/N-weighted stack and a ‘boxcar’ method to specifically isolate the high-velocity outflow component. In contrast, Glazer et al. (2025) measured the velocity centroid of a composite LIS profile, created by averaging five transitions (Si ii $\lambda$1260, O i $\lambda$1302, Si ii $\lambda$1304, C ii $\lambda$1334, and Si ii $\lambda$1526), which reflects the bulk motion of the total cool gas reservoir. Thus, the lower velocity found by Glazer et al. (2025) is a measure of the average gas kinematics and does not contradict our finding of a faster, distinct outflow component. Future work applying a consistent methodology to both datasets will be valuable to harmonize these complementary views of gas kinematics in the early universe.
Combining the results from Figures 4, 5, and 8, this lack of significant evolution in the cool outflow properties at $z>3$ provides a powerful test for theoretical models of galaxy feedback in the early universe. The scaling relations of outflow velocity and equivalent width with stellar mass at $z\sim 3-6$ are roughly consistent with those observed at $z\sim 2$. This suggests that a persistent, unevolving feedback mechanism is already in place at $z>3$. The strong outflows we detect indicate that stellar feedback remains efficient at driving gas out of galaxies even in the compact, dense environments typical of high-redshift systems. Our findings support a scenario where the interplay between star formation and feedback follows the same physical laws from cosmic noon out to $z\sim 6$, without requiring a fundamental transition in the mode of feedback for the population probed by our sample.